Y=x-4
Y=4x-10
Solve the system of linear equations using substitution.

Answer by algebrahouse.com(1659)

y = x - 4
y = 4x - 10

x - 4 = 4x - 10 {substituted (x - 4), in for y, into bottom equation}
-4 = 3x - 10 {subtracted x from each side}
6 = 3x {added 10 to each side}
x = 2 {divided each side by 3}

y = x - 4 {top original equation}
y = 2 - 4 {substituted 2, in for x, into top original equation}
y = -2 {subtracted}

x = 2 and y = -2

To ask a question, visit: www.algebrahouse.com